a body travels 2 m in 2nd second & 6 m in next 4 second. what will be the distance travelled in 9th second?

Grade:9

1 Answers

A M S ARUN KRISHNA
8 years ago
9 metres
v =d / t
v = 2 / 2 = 1
d = vt
d = 1 × 9 = 9 metres